A boy saves part of his allowance by placing in a piggy-bank 25 centavo coins and 50 centavo coins. If after two months he finds out that his total saving is 130 pesos and the number of fifty centavos is forty less than twice that of the twenty-five centavo coins, how many of each coin has he saved?
Answer by TimothyLamb(4379)

---
x = 25 cent coins
y = 50 cent coins
---
25x + 50y = 13000
y = 2x - 40
---
put the system of linear equations into standard form
---
25x + 50y = 13000
2x - y = 40
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = 25 cent coins = 120
y = 50 cent coins = 200
---
system matrix
25 50
2 -1
---
inverse of system matrix
0.008 0.4
0.016 -0.2
---
determinant of system matrix = -125
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 4-equations 4-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php